OpenStudy (anonymous):

what is the least common denominator of 4/9 and 8/15

OpenStudy (anonymous):

it will be 45

OpenStudy (anonymous):

how lets see that 9 = 3*3 15 = 3*5 3 is common in both so, 3*3*5=9*5=45